The supply of corn by the United States is Qa = a + bp, and the supply by the rest of the world is Qr = a + ep. What is the total world supply?
Answer #1

Answer : Given,

Supply of U.S. : Qa = a + bp

Supply of rest world : Qr = a + ep

Now,

Let, world supply = Qw. So,

Qw = Qa + Qr = (a + bp) + (a + ep)

=> Qw = a + bp + a + ep

=> Qw = 2a + bp + ep

=> Qw = 2a + p (b + e)

Therefore, the total world supply is, Qw = 2a + p (b + e).